OpenStudy (anonymous):

Find the logarithm log(1/4) (1/64).

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

can you retype the question using the Equation editor please?

OpenStudy (anonymous):

\[\log_{\frac{ 1 }{ 4}} (\frac{ 1 }{ 64 })\]

OpenStudy (anonymous):

\[\log_{1/4} (1/64)\]

OpenStudy (anonymous):

Ranga i got to go now Can you leave an example n how to do it an dill figure it out please

OpenStudy (ranga):

ok

OpenStudy (ranga):

IF \[\Large \log_{B} A = C\]THEN\[\Large A = B^{C}\]Let\[\Large \log_{\frac{ 1 }{ 4 }} \frac{ 1 }{ 64 } = x\]then\[\Large \frac{ 1 }{ 64 } = (\frac{ 1 }{ 4 })^{x} = \frac{ 1^{x} }{ 4^{x} } = \frac{ 1 }{ 4^{x} }\]\[\Large 4^{x} = 64 = 4^{3}\]So x = 3

OpenStudy (anonymous):

This is History, not math. Please post it in the Mathematics group. This is considered spam and is against the Code of Conduct which OpenStudy follows. You can read the Code of Conduct at http://OpenStudy.com/code-of-conduct . Thanks